Bonferroni's method ensures that the family-wise error rate is less than or equal to the design level such as 0.05 after all possible pairs-wise comparisons have been made. Three groups have 3 possible pairs-wise comparisons. Four groups have 6 possible pairs-wise comparisons.

Please answer following questions 26-29 based on this method.

A study has three groups. Given an overall design error rate 0.05, a pairs-wise comparison can be claimed statistically significant difference when the p-value less than

Group of answer choices

0.05

0.017

0.15

0.008

0.3

Flag this Question

Question 273.33 pts

A study has four groups. Given an overall design error rate 0.05, a pairs-wise comparison can be claimed statistically significant difference when the p-value less than

Group of answer choices

0.05

0.017

0.15

0.008

0.3

Flag this Question

Question 283.33 pts

A study has three groups. An investigator wants to claim a statistically significant difference if a pairs-wise comparison p-value less 0.05. The overall design error rate for this study could be as large as

Group of answer choices

0.05

0.017

0.15

0.008

0.3

Flag this Question

Question 293.33 pts

A study has four groups. An investigator wants to claim a statistically significant difference if a pairs-wise comparison p-value less 0.05. The overall design error rate for this study could be as large as

Group of answer choices

0.05

0.017

0.15

0.008

0.3
Answer #1

A study has three groups. Given an overall design error rate 0.05, a pairs-wise comparison can be claimed statistically significant difference when the p-value less than 0.05/3=0.017

Option: 0.017

A study has four groups. Given an overall design error rate 0.05, a pairs-wise comparison can be claimed statistically significant difference when the p-value less than 0.05/6=0.008

Option: 0.008

A study has three groups. An investigator wants to claim a statistically significant difference if a pairs-wise comparison p-value less 0.05. The overall design error rate for this study could be as large as 3*0.05=0.15.

Option: 0.15

A study has four groups. An investigator wants to claim a statistically significant difference if a pairs-wise comparison p-value less 0.05. The overall design error rate for this study could be as large as 6*0.05=0.30.

Option: 0.3